OpenStudy (anonymous):

Which is the solution of the system? {x+y=-2 {2x-3=-9 A. (-4, 2) B. (-3, 1) C. (-3, -5) D. (4, -6)

OpenStudy (diyadiya):

2x-3=-9 Add 3 on both sides can you do that?

OpenStudy (anonymous):

nope i didn't but i just need the answr tryna finish this test asap

OpenStudy (diyadiya):

Sorry! We can't give you "just answers" but if you want i can help you to find it :) Please Read the Code of Conduct > http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

i know but so i just add the 3 and it' ll give me the answer?

OpenStudy (anonymous):

too both of the sides?

OpenStudy (diyadiya):

No that's just the first step 2x-3=-9 2x-3 + 3 = -9 +3 2x =-9 + 3 2x =-6 now divide both sides by 2 2x/2 = -6/2 x = -3 Now you have to substitute the value of x (that is -3 ) in the equation x+y=-2 can you try that?

OpenStudy (anonymous):

yyeah i got it

OpenStudy (diyadiya):

:)

OpenStudy (anonymous):

thanks

OpenStudy (diyadiya):

Anytime :)

OpenStudy (anonymous):

Can u help me out with this Which of the following relations is a function? *A. {(2, 2), (3, 3), (4, 4), (5, 5)} *B. {(2, 1), (2, 3), (3, 6), (3, 7)} *C. {(3, 5), (4, 9), (4, 11), (5, 14)} *D. {(4, 8), (6, 6), (8, 4), (8, 2)}